# Safeguarding Mobile GUI Agent via Logic-based Action Verification

### v1

@@ 8.2. Verification Accuracy
Figure 5 presents the verification accuracy of different verification methods, evaluated using standard metrics: Accuracy, Precision, Recall, and F1 Score. Notably, VSA consistently outperforms reflection-based methods across all metrics and datasets, achieving near-perfect accuracy on the LlamaTouch dataset. One interesting observation is the high recall performance of Pre-action Reflection on the Challenge Dataset. However, this comes at the cost of a high false positive rate (60%), leading to excessive error flagging when instructions become complex. As a result, while it has a higher likelihood of detecting any existing error (high Recall), the probability that a flagged error is actually an error remains low (low Precision).
@@
Analysis. For the relatively simpler LlamaTouch dataset, Post-action Reflection achieves performance comparable to VSA. This is because it leverages the action’s resulting screen state as additional context for verification, boosting accuracy compared to Pre-action Reflection. However, it has critical limitations: it cannot correct irreversible actions. Moreover, our results indicate that this advantage diminishes significantly, as instructions become more complex and involve longer action sequences (as seen in the Challenge Dataset). This trend highlights the compounding effect of errors stemming from over-reliance on LFMs for verification.
@@
In contrast, VSA enables pre-action verification while maintaining high performance across both datasets. This demonstrates the effectiveness and scalability of VSA’s design, which encodes the specification once and performs subsequent verification in a rule-based manner, regardless of task complexity or length. Notably, VSA-Warm achieves nearly perfect results on the LlamaTouch dataset and significantly surpassing all other methods on the Challenging dataset. This demonstrates the potential of VSA’s design when equipped with a well-constructed memory of candidate state predicates. While the construction of such a memory (e.g., offline exploration, user demonstrations, runtime analysis) is beyond the scope of this paper (see § 10), the results clearly underscore the promise of VSA.
Randomness Mitigation.
@@
Although VSA employs only one LFM query per task, when repeatedly verifies the same instruction, inconsistent results may still arise due to the probabilistic nature of LFM. The same incosistency applies to all methods that includes LFM in its process, including reflection methods. To mitigate this randomness, a majority voting approach can be employed, wherein the verification process is repeated $N$ times, and the final outcome is determined by majority voting (referred to as Majority at N).
Figure 6 illustrates the each method’s verification accuracy for 50 instructions in the Challenge dataset as $N$ increases. We observed that while reflection-based schemes consistently shows lower accuracy, both VSA-Cold and VSA-Warm show significant improvements, reaching 96% and 100%, respectively. This disparity stems from differences in their verification mechanisms: reflection schemes produce direct true/false responses with lower randomness, whereas both VSA-Cold and VSA-Warm generate logical formulas with multiple predicates through the LFM, making them more susceptible to stochastic variations. These findings suggest that VSA has significant room for performance enhancement through majority voting for verification. However, it is worth noting that excessively increasing the number of verification iterations can introduce considerable overhead in terms of latency and cost. A detailed analysis of this overhead issue is provided in § 8.4.
@@ 8.3. Effectiveness of Feedback Generation
The usefulness of a verification system lies not just in identifying errors, but in improving the GUI agent’s performance through feedback generation. This section compares the effectiveness of VSA’s logic-based feedback generation with that of the LFM-based feedback used in reflection methods.
To evaluate this, we instructed the M3A GUI agent to execute 25 challenging instructions from our Challenge dataset. Initially, without external feedback (Agent w/o feedback), M3A successfully completed only 10 out of 25 tasks. Next, we integrated each verification system to M3A and re-executed each instruction, observing whether previously failed tasks could now be correctly performed. Note that we evaluated using our own dataset over LlamaTouch dataset because LlamaTouch’s ground truth fails to account the multiple paths often present in real-world mobile tasks. On the other hand, our Challenge dataset includes hand-crafted ground truth encompassing multiple valid execution paths, providing more accurate evaluation than LlamaTouch’s single-path approach which showed poor correlation with human evaluation ($<$60% fidelity).
Figure 7 summarizes these results. Notably, neither pre-action nor post-action reflection successfully corrected any previously failed tasks. In contrast, VSA-Cold feedback corrected 9 out of 15 tasks (60%), and VSA-Warm successfully corrected 13 tasks (86.7%). It is worth noting that VSA-Warm’s two failures stemmed from unconventional application designs, not from the inability to understand the execution context. For instance, a clock application required time input as a single numerical sequence (hour, minute, second combined) rather than separate components. This unconventional interaction hinders VSA ’s ability to provide meaningful feedback. Nevertheless, given that M3A is a state-of-the-art GUI agents powered by one of the most powerful LFM (gpt-4o), achieving this significant enhancement underscores VSA’s effectiveness.
This superior performance stems from VSA’s ability to generate precise, deterministic feedback, clearly indicating what aspects of the agent’s execution need to be corrected. Furthermore, VSA’s roadmap feedback proactively guides the GUI agent toward the correct execution path, preventing errors from occurring at the first place. For example, with the instruction: ”Go to the r/technology subreddit. In the most controversial post in your field over the past month, go to the comments on that post. Find the most controversial comment in the comments, and upvote that comment,” the M3A agent w/o feedback repeatedly toggled the upvote action due to unclear state tracking, not knowing that it has already upvoted. Yet, when given a clear guideline saying that it has fulfilled the precondition to set ‘upvote’ to ‘true’, M3A to successfully complete the task without unnecessary repetition.
In contrast, reflection methods, which rely on an LFM for feedback generation, frequently produce misleading guidance and fails to correct any of the previously failed instructions. Furthermore, we observed cases where reflection initially flagged an action as incorrect but reversed its decision after the GUI agent repeated the same action, illustrating inherent indecisiveness and unreliability in its feedback generation.
One potential limitation of VSA, though not observed in our evaluation, is that false positive errors could prematurely terminate tasks that the GUI agent might have otherwise completed successfully. Unlike reflection-based methods, where an agent can typically recover quickly– within 1~3 actions – after a misleading feedback, a single incorrect predicate in the Specification could cause VSA to halt task execution entirely. Nevertheless, our evaluation clearly demonstrates that the practical benefits of VSA’s decisive feedback significantly outweigh the potential risks of rare false positives. Particularly in safety-critical systems, such as automated mobile tasks, strictly preventing erroneous actions—even at the expense of occasional false positives—is significantly more beneficial than allowing potentially harmful actions to proceed, as traditional reflection methods currently do.
@@ 9. Related Work
Validating LFM responses. The probabilistic nature of LFMs often leads to hallucinated or inaccurate content. Numerous methods have been proposed to mitigate these issues, including Chain-of-Verification (CoVe) (Dhuliawala et al., 2023), concept-level validation and rectification (EVER) (Kang et al., 2023), self-refinement  (Madaan et al., 2023), and retrieval-augmented knowledge grounding (Re-KGR)  (Niu et al., 2024). These approaches typically leverage additional LLM invocations for self-reflections (Li et al., 2024a; Wang et al., 2024) or rely on external validation against ground-truth data or knowledge bases  (Min et al., 2023; Chern et al., 2023; Nan et al., 2021; Goodrich et al., 2019; Shuster et al., 2021). However, in mobile task automation, actions of a mobile GUI agent lack a ground truth and its semantic ambiguity makes verification exceptionally difficult. VSA effectively handles this challenge by designing a domain-specific language (DSL, § 4.1) and developer library (§ 4.2) that effectively capture the underlying semantics of these ambiguous GUI actions.
Leveraging Formal Logic for AI Safety. Recent research has explored the use of formal logic to ensure the safety and reliability of AI agents, particularly in areas like planning and control (Liu et al., 2023; Yang et al., 2024; Wu et al., 2024; Obi et al., 2025; Fuggitti and Chakraborti, 2023; Cosler et al., 2023). Notably, SELP (Wu et al., 2024) ensures safe and efficient robot task planning by integrating constrained decoding and domain-specific fine-tuning. Ziyi et al. (Yang et al., 2024) proposed a safety module for LLM-based robot agents by converting natural language safety constraints into Linear Temporal Logic. However, while previous works typically focus on agents with a pre-defined safety concerns such as map navigation or indoor task planning  (Liu et al., 2023; Yang et al., 2024; Wu et al., 2024; Obi et al., 2025), our system targets agents with a enormous number of environment, i.e., mobile apps. To handle different environments of mobile apps, our optimized autoformalization algorithm effectively translates safety requirements for diverse set of mobile tasks.
@@ 10. Discussion & Future Work
Piggybacking on App Development Frameworks. The current VSA implementation relies on developer-defined state and transition definition. However, given that modern mobile development frameworks like React Native, Jetpack Compose, and SwiftUI (Meta, 2025; Google, 2025; Apple, 2025) already incorporate built-in state managements, there is significant potential to ”piggyback” on these existing mechanisms. Integrating VSA in this way could streamline the development process and reduce the overhead associated with adopting VSA.
Automated State Definition via LFM. Another approach for simplifying VSA’s integration is automating the state definition process using LFMs. Existing GUI agents (Wen et al., 2023; Lee et al., 2024; Zhang et al., 2023) often employ LLM-powered screen analysis to automatically annotate GUI elements and infer application state. A similar approach could be adapted to identify and define candidate state predicates for VSA. However, while this automation could reduce developer effort, it introduces inherent risks due to the probabilistic nature of LLMs. Relying on potentially inaccurate LLM-derived state definitions could compromise the system’s reliability. Therefore, while LLM-based automation offers a potential convenience, the current approach of developer-defined states ensures the precision necessary for a robust verification system.
@@ 11. Conclusion
This paper introduced VeriSafe Agent (VSA), a novel logic-based verification system designed to enhance the safety and reliability of Mobile GUI Agents. Through VSA, we have demonstrated that logic-driven verification can effectively safeguard GUI agents against the inherent uncertainties of LFM-based task automation. As AI-driven tasks becomes increasingly integrated into our daily life, VSA presents a crucial step toward safer use of AIs.

### v2

@@ 8.2. Verification Accuracy
Figure 5 presents the verification accuracy of different verification methods, evaluated using standard metrics: Accuracy, Precision, Recall, and F1 Score. Notably, VSA consistently outperforms reflection-based methods across all metrics and datasets, achieving near-perfect accuracy on the LlamaTouch dataset.
@@
For simple tasks in LlamaTouch dataset, post-action reflection achieves significantly higher accuracy compared to pre-action reflection. This is because it leverages the action’s resulting screen as additional context for verification. However, it has critical limitations: it cannot correct irreversible actions. Moreover, our results indicate that this advantage diminishes significantly as instructions become more complex and involve longer action sequences (as seen in the Challenge dataset). This trend highlights the compounding effect of errors stemming from over-reliance on LLMs for verification.
Error Type Variant LlamaTouch Challenge
Overall Missing Predicate VSA-Cold 76/250 24/50
VSA-Warm 24/250 14/50
Critical Missing Predicate VSA-Cold 10/250 3/50
VSA-Warm 3/250 2/50
Superfluous Predicate VSA-Cold 7/250 4/50
VSA-Warm 0/250 0/50
Constraint Mismatch VSA-Cold 8/250 11/50
VSA-Warm 4/250 6/50
Table 1. VSA predicate translation accuracy
@@
In contrast, VSA performs pre-action verification while maintaining high performance across both datasets, thanks to its design that encodes the specification once and performs subsequent verification in a rule-based manner, regardless of task complexity or length. Notably, VSA-Warm achieves near-perfect results on LlamaTouch and significantly outperforms all baselines on Challenge, highlighting the benefit of a well-constructed predicate memory. While building such a memory (e.g., via offline exploration, user demonstrations, runtime analysis) is beyond this paper’s scope (§ 10), these results underscore its promise.
An interesting observation is the high recall of Pre-action Reflection on the Challenge Dataset. However, this comes at the cost of a high false positive rate (72%), leading to excessive error flagging when instructions become complex. While this increases the chance of detecting any error, the likelihood that a flagged action is truly erroneous remains low (low precision).
Predicate Translation Accuracy. To gain deeper insight into the verification results, we further evaluated VSA at the state predicate level. We manually constructed a ground-truth set of predicates $P_{GT}$ for each instruction and compared it against $P_{Spec}$ generated by VSA.
VSA’s Verification errors fall into three categories: 1) Missing Predicate — necessary predicates absent from the specification, potentially allowing incorrect behaviors (False Negatives); 2) Superfluous Predicate — unnecessary predicates included, potentially rejecting correct behaviors (False Positives); 3) Constraint Mismatch — correct predicate but with inaccurate constraints, which can cause both False Negatives and False Positives.
Table 1 shows the counts of such errors. Note that Overall Missing Predicate counts all predicates absent relative to $P_{GT}$, which defines the maximal set of possible predicates (including semantically overlapping ones). In practice, VSA tends to select a minimal set of core constraints, which can lead to an overestimation of “missing” predicates. We therefore also report Critical Missing Predicates, which are those whose absence directly affects verification results.
A key observation is that VSA-Warm significantly reduces both missing and superfluous predicates. This demonstrates that predicate memory effectively improves translation accuracy by narrowing the search space for necessary predicates. Furthermore, the distribution of errors across all three types indicates that, despite VSA’s strong overall verification performance, residual risks from predicate translation remain. In the Discussion section, we propose complementary policies to mitigate these risks and further enhance VSA’s safety.
Randomness Mitigation. Although VSA issues only a single LFM query per task, repeated verification of the same instruction may yield inconsistent results due to the stochasticity of LFMs. The same inconsistency applies to all methods that invoke an LFM, including reflection-based approaches. To mitigate this randomness, we employ a majority-voting scheme that repeats verification $N$ times and returns the majority decision (Majority-at-$N$).
Figure 6 reports accuracy on 50 instructions from the Challenge dataset as $N$ increases. Reflection-based schemes show limited gains and remain less accurate overall, whereas both VSA-Cold and VSA-Warm improve substantially, reaching 96% and 100%, respectively.
This performance disparity stems from differences in verification mechanisms: reflection schemes produce direct true/false judgments with relatively low variance, whereas VSA encodes a logical specification with multiple predicates—small stochastic differences can flip individual predicate decisions—making majority voting especially effective at averaging out noise.
These findings indicate that VSA can achieve substantial performance gains through majority voting. However, increasing verification iterations introduces considerable overhead in terms of latency and cost. We analyze this trade-off in detail in § 8.4.
@@ 8.3. Effectiveness of Feedback Generation
The usefulness of a verification system lies not just in identifying errors, but in improving the GUI agent’s performance through feedback generation. This section evaluates VSA’s logic-based feedback generation against the LLM-based feedback used in reflection methods.
We instructed the M3A GUI agent to execute 25 challenging instructions from our Challenge dataset. Initially, without external feedback, M3A successfully completed only 10 of 25 tasks (40% success rate). We then integrated each verification system with M3A and re-executed the failed instructions to observe whether the feedback can correct the previously failed tasks.
Figure 7 summarizes the results. Notably, Neither pre-action nor post-action reflection corrected any previously failed tasks. In contrast, VSA-Cold corrected 9 of 15 failures (60%), and VSA-Warm corrected 13 of 15 (86%). The two remaining failures under VSA-Warm were due to unconventional app designs rather than misunderstanding of context. For instance, one clock application required time input as a single numerical sequence (e.g., ”143000” for 14:30:00) rather than separate hour, minute, and second components. Such unconventional interfaces could limit VSA’s ability to provide meaningful feedback. Nevertheless, given that M3A employs GPT-4o—one of the most capable LLMs—achieving this performance improvement through systematic harness demonstrates VSA’s effectiveness.
VSA’s superior performance stems from its ability to generate precise, deterministic feedback that clearly identifies what needs correction (i.e., unmet predicates). Moreover, VSA’s roadmap feedback proactively guides agents toward correct execution paths, preventing errors from happening in the first place. For example, when asked to upvote the most controversial comment in a subreddit, M3A without feedback repeatedly toggled the upvote button because it could not identify whether the upvote has been properly set. However, when VSA provided clear guidance that the precondition to set ’upvote’ to ’true’ had been satisfied, M3A completed the task without unnecessary repetition.
@@
By contrast, reflection methods rely on LFM-generated narrative feedback that is often misleading or inconsistent; we observed cases where a reflection method initially flagged an action as incorrect but reversed its judgment when the agent repeated the same action, indicating indecisive and unreliable guidance.
A potential limitation of VSA, though not observed in our evaluation, is that false positives could prematurely terminate tasks that the agent might otherwise complete successfully. Unlike reflection-based methods, where agents typically recover within 1–3 actions after misleading feedback, a single incorrect predicate in VSA’s specification could halt task execution entirely. However, our evaluation demonstrates that VSA’s decisive feedback provides practical benefits that significantly outweigh the risks of rare false positives. In safety-critical applications such as automated mobile tasks, preventing erroneous actions—even at the cost of occasional false positives—is preferable to allowing potentially harmful actions to proceed.
@@ 8.4. Latency and Cost
8.5. Developer Burden
To estimate the developer effort required for VSA integration, we analyzed 18 commercial mobile applications from our dataset. Since these applications are closed-source, we simulated the integration process by manually enumerating the verification-relevant state predicates for each of the 18 target apps. We then estimated the lines of code (LoC) by multiplying predicate counts by the average LoC per predicate measured from our reference implementation.
On average, each application required 58.77 predicates, translating to approximately $\sim$437 LoC. Across all 18 apps, we generated 1,058 predicates in total, of which 540 (51%) were activated during our experiments. The number of predicates varied with application complexity. Google Tasks, a single-function app, required only 22 predicates, while feature-rich applications like Instagram, X (Twitter), and Facebook required 81, 91, and 72 predicates, respectively. This correlation between application complexity and predicate count is expected, as more complex applications naturally have more states to verify.
While these estimates suggest a non-negligible implementation effort, several factors mitigate the burden. First, many predicates follow standard patterns that could be templated or automated. Second, the predicate definitions align naturally with existing application state logic that developers already maintain. Third, predicate granularity provides a practical lever for balancing coverage and effort, allowing developers to selectively implement only safety-critical predicates.
Overall, the effort required for predicate definition represents a one-time investment that yields ongoing benefits in error prevention and agent reliability. For applications where automated agents perform high-stakes tasks, the safety improvements demonstrated in our experiments justify this initial development effort.
@@ 9. Related Work
Validating LFM responses. The probabilistic nature of LFMs often leads to hallucinated or inaccurate content. Numerous methods have been proposed to mitigate these issues, including Chain-of-Verification (CoVe) (Dhuliawala et al., 2023), concept-level validation and rectification (EVER) (Kang et al., 2023), self-refinement  (Madaan et al., 2023), and retrieval-augmented knowledge grounding (Re-KGR)  (Niu et al., 2024). These approaches typically leverage additional LFM invocations for self-reflections (Li et al., 2024a; Wang et al., 2024) or rely on external validation against ground-truth data or knowledge bases  (Min et al., 2023; Chern et al., 2023; Nan et al., 2021; Goodrich et al., 2019; Shuster et al., 2021). However, in mobile task automation, actions of a mobile GUI agent lack a ground truth and its semantic ambiguity makes verification exceptionally difficult. VSA effectively handles this challenge by designing a domain-specific language (DSL, § 4.1) and developer library (§ 4.2) that effectively capture the underlying semantics of these ambiguous GUI actions.
Leveraging Formal Logic for AI Safety. Recent research has explored the use of formal logic to ensure the safety and reliability of AI agents, particularly in areas like planning and control (Liu et al., 2023; Yang et al., 2024; Wu et al., 2024; Obi et al., 2025; Fuggitti and Chakraborti, 2023; Cosler et al., 2023). Notably, SELP (Wu et al., 2024) ensures safe and efficient robot task planning by integrating constrained decoding and domain-specific fine-tuning. Ziyi et al. (Yang et al., 2024) proposed a safety module for LFM-based robot agents by converting natural language safety constraints into Linear Temporal Logic. However, while previous works typically focus on agents with a pre-defined safety concerns such as map navigation or indoor task planning  (Liu et al., 2023; Yang et al., 2024; Wu et al., 2024; Obi et al., 2025), our system targets agents with an enormous number of environment, i.e., mobile apps. To handle different environments of mobile apps, our optimized autoformalization algorithm effectively translates safety requirements for diverse set of mobile tasks.
@@ 10. Discussion & Future Work
Piggybacking on App Development Frameworks. The current VSA implementation relies on developer-defined state and transition definition. However, given that modern mobile development frameworks like React Native, Jetpack Compose, and SwiftUI (Meta, 2025; Google, 2025; Apple, 2025) already incorporate built-in state management, there is significant potential to ”piggyback” on these existing mechanisms. Integrating VSA in this way could streamline the development process and reduce the overhead associated with adopting VSA.
Automated State Definition via LFM. Another approach for simplifying VSA’s integration is automating the state definition process using LFMs. Existing GUI agents (Wen et al., 2023; Lee et al., 2024; Zhang et al., 2023) often employ LFM-powered screen analysis to automatically annotate GUI elements and infer application state. A similar approach could be adapted to identify and define candidate state predicates for VSA. However, while this automation could reduce developer effort, it introduces inherent risks due to the probabilistic nature of LFMs. Relying on potentially inaccurate LFM-derived state definitions could compromise the system’s reliability. Therefore, while LFM-based automation offers a potential convenience, the current approach of developer-defined states ensures the precision necessary for a robust verification system.
@@
Interactive Feedback & Dual Verification While VSA achieves high verification accuracy, it cannot guarantee 100% safety. This is critical in high-risk scenarios like financial transactions where even a single error is unacceptable. To address these limitations, we propose human-in-the-loop strategies. Specifically, we recommend two complementary safeguards: (1) interactive rule-level confirmation that prompts users for confirmation when rules are repeatedly violated, enabling quick human identification of potential False Positives. (2) Dual verification requiring human approval for critical operations (payments, data transfers) regardless of VSA’s verification outcome—which guards against potential False Negatives. These mechanisms provide practical safeguards for commercial deployment while preserving VSA’s automation benefits.
@@ 11. Conclusion
This paper introduced VeriSafe Agent (VSA), a novel logic-based verification system designed to enhance the safety and reliability of Mobile GUI Agents. Through VSA, we have demonstrated that logic-driven verification can effectively safeguard GUI agents against the inherent uncertainties of LFM-based task automation. As AI-driven tasks become increasingly integrated into our daily life, VSA presents a crucial step toward safer use of AIs.
Acknowledgements.
This work was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant (RS-2019-II190421, RS-2022-II221045, RS-2025-25442569, RS-2025-02305705, RS-2020-II201819, RS-2023-00232728), the National Research Foundation of Korea(NRF) grant (RS-2024-00338454, RS-2021-NR060080, RS-2025-00522352, RS-2024-00347516), and K-Startup grant (20144069) funded by Korea government (MSIT).